# Snail-Radar: A large-scale diverse dataset for the evaluation of 4D-radar-based SLAM systems

### v1

@@ 9 Conclusions
This paper presents a large-scale 4D radar dataset designed for localization and mapping applications based on multi-sensor fusion. The dataset was collected using three platforms under diverse environmental conditions, including rainy days, nights, campus roads, and tunnels. A total of 44 sequences were repeatedly collected over eight routes.
Reference poses (at the start and end for large-scale sequences) are generated using a LIO method that sequentially localizes the undistorted Hesai lidar frames to the TLS map. A data reversion technique enables backward LIO in time. Sensor data are synchronized using a two-step scheme which first uses the hardware-synced lidar data to map the message host times to GNSS times to remove jitter and long-term drift, and secondly uses odometry and correlation algorithms to estimate the constant time offsets between sensor data and the GNSS/INS solution. The extrinsic parameters between all sensors are refined using several methods and are provided with the dataset. We hope our dataset will facilitate research into the newly advented 4D radars, particularly in odometry, mapping, and place recognition.
@@
We acknowledge the assistance of Xiaochen Wang and Hanwen Qi in collecting the TLS data with the RTC360, lent by Professor Xinlian Liang. We are also grateful to Shangjia Liu and Qi Liu for their help with data collection.

### v3

@@ 10 Conclusions
This paper introduces a large-scale 4D radar dataset designed for localization and mapping applications leveraging multi-sensor fusion. The dataset was collected using three platforms across diverse environmental conditions, including rainy days, nighttime, campus roads, and tunnels. In total, 44 sequences were repeatedly recorded over eight distinct routes.
Accurate reference poses (at the start and end of large-scale sequences) were generated using a LIO method that sequentially localizes undistorted Hesai lidar frames to a TLS map. A data inversion technique enables backward LIO processing, which in turn allows two-way reference trajectory smoothing and quality assessment of the resulting reference trajectories.
Sensor data synchronization was achieved using a two-step scheme. First, hardware-synchronized lidar data was used to map message host times to GNSS times, eliminating jitter and long-term drift. Second, odometry and correlation algorithms were employed to estimate constant time offsets between the sensor data and the GNSS/INS solution.
Using this dataset, we evaluated several recent radar-based odometry and place recognition methods, showing the challenges inherent to 4D radar data. Additionally, as the dataset provides dense reference point clouds, it is well-suited for evaluating neural 3D reconstruction methods (Ming et al., 2025).
@@
We acknowledge the assistance of Xiaochen Wang and Hanwen Qi in collecting the TLS data with the RTC360, lent by Prof. Xinlian Liang. We are also grateful to Shangjia Liu and Qi Liu for their help with data collection, to Dr. Yusheng Wang, Jian Zhou, and Hailiang Tang for fruitful discussions, to Prof. Charles Toth and the anonymous reviewers for their helpful and stimulating suggestions.
{funding}
This research is partly funded by the National Key Research and Development Program of China (International Scientific and Technological Cooperation Program) under Grant 2022YFE0139300, the National Natural Science Foundation of China under Grant 42374047, and the Fundamental Research Fund Program of LIESMARS.